OpenStudy (anonymous):

A cylindrical vase has a diameter of 6 inches. At the bottom of the vase, there are 8 marbles, each of diameter 3 inches. The vase is filled with water up to a height of 10 inches. Which of the following could be used to calculate the volume of water in the vase?

OpenStudy (anonymous):

can someone please help me?

OpenStudy (anonymous):

@AllTehMaffs

OpenStudy (anonymous):

@emilybbydoll

OpenStudy (anonymous):

@jango_IN_DTOWN

OpenStudy (dayakar):

do u know how to find volume of cylinder

OpenStudy (anonymous):

A. π(2ft)2(10ft) + one over threeπ(13ft − 10ft)2(2ft) B. π(10ft)2(2ft) + one over threeπ(13ft − 10ft)2(2ft) C. π(2ft)2(10ft) + one over threeπ(2ft)2(13ft − 10ft) D.π(10ft)2(2ft) + one over threeπ(2ft)2(13ft − 10ft)

OpenStudy (anonymous):

yes with πr^2h

OpenStudy (jango_in_dtown):

And each marble is a sphere. Volume of a sphere is?

OpenStudy (anonymous):

4/3πr^3?

OpenStudy (jango_in_dtown):

correct

OpenStudy (dayakar):

find volume of a vase using above formula

OpenStudy (anonymous):

o I'm sorry those are not the answers:(

OpenStudy (anonymous):

A. π(10in)2(3in) − 1.5(four over threeπ(8in)3) B. π(3in)2(10in) − 1.5(four over threeπ(8in)3) C. π(10in)2(3in) − 8(four over threeπ(1.5in)3) D. π(3in)2(10in) − 8(four over threeπ(1.5in)3)

OpenStudy (anonymous):

The volume equals 282.6

OpenStudy (dayakar):

volume of water in the vase = volume of cyliderical vase - 8* spehere volume

OpenStudy (anonymous):

how would I find the sphere volume?

OpenStudy (jango_in_dtown):

diameter=3 so radius =1.5

OpenStudy (anonymous):

14.13?

OpenStudy (dayakar):

\[spehere volume = 4/3 * \pi *r^{3}\]

OpenStudy (anonymous):

14.14

OpenStudy (dayakar):

\[\pi *3 ^{2}*10 -8*\frac{ 4 }{ 3 }* \pi*(1.5)^3\]

OpenStudy (dayakar):

do u understand

OpenStudy (anonymous):

yes thank you so much

OpenStudy (anonymous):

its was right!:)